OpenStudy (anonymous):

Last year in Kerry’s school, the ratio of students to teachers was 22:1. How many students were there if there were 8 teachers? A)184 B)168 C)176 D)192

OpenStudy (anonymous):

Let no. of students be x, teachers y Ratio = x/y = 22/1 No. of teachers given is 8\[\frac{ x }{ 8}=\frac{ 22 }{ 1 }\] Cross multiply and solve for x. The result is the number of students.

OpenStudy (anonymous):

C)?

OpenStudy (texaschic101):

yep...it is C

OpenStudy (anonymous):

ok thanks :)

OpenStudy (texaschic101):

don't thank me....thank d3Xter for that great explanation.